OpenStudy (anonymous):

\[\int\limits \frac{dx}{x \sqrt{81x ^{2}-16}}\]

OpenStudy (anonymous):

Trig Sub! I know how to do these now! :-D One moment, I'll show you steps... I need to get my lunch from the microwave oven :-3

OpenStudy (anonymous):

First while I go to get this food, 81 = 9^2 and 16 = 4^2

OpenStudy (anonymous):

Trig sub. For this form with the 1/(x*sqrt(b^2*x^2-a^a)).... let \(x = \large\frac{4}{9}\sec\theta\) \(dx = \large\frac{4}{9}\tan\theta\sec\theta\) @Beatles watch what happens when you put this in place for "x" and "dx" :-)

OpenStudy (anonymous):

\[\int\limits \frac{1}{x\cdot (9^2x^2-4^2)^{1/2}}\cdot dx\] \[\int\limits \frac{1}{x\cdot (9^2(\sec\theta)^2-4^2)^{1/2}}\cdot (\tan\theta\sec\theta \ \ d\theta)\]

OpenStudy (anonymous):

Take the unit circle identity for sine and cosine, divide both side by cos^2 \[\frac{\sin^2\theta}{\cos^2\theta} + \frac{\cos^2\theta}{\cos^2\theta} = \frac{1}{\cos^2\theta}\] Can you see what happens here, @Beatles? Can you write the identity here in terms of tangent-squared now?

OpenStudy (anonymous):

Still there m8? :-D

OpenStudy (experimentx):

lol ... the fastest way ... do opposite.

OpenStudy (anonymous):

can we use the theorem of inverse trigo func. -_-

OpenStudy (anonymous):

\[\int\limits \frac{1}{\sec\theta \cdot (9^2(\sec\theta)^2-4^2)^{1/2}}\cdot (\tan\theta\sec\theta \ \ d\theta)\] @experimentX actually faster way is to use the integral tables ;-)

OpenStudy (anonymous):

Err I mean wait that's 4/9 sec theta

OpenStudy (anonymous):

Let me fix that...

OpenStudy (experimentx):